# The Quantum-Cryptographic Co-evolution

### 03 Apr 2026 v1

@@ Abstract
As quantum computing matures toward the realization of Cryptographically Relevant Quantum Computers (CRQC), global cryptographic infrastructure faces an existential threat. This paper introduces a two-dimensional coordinate system to map the co-evolution of cryptographic resilience ($x$-axis) and computational capability ($y$-axis). By analyzing the four resulting quadrants, we categorize the transition from legacy classical systems to quantum-resilient architectures. We argue that the ”Quantum Gap”—the delta between CRQC arrival and quantum-safe adoption—represents the highest systemic risk, necessitating an immediate transition to crypto-agile frameworks.
@@ I Introduction
With the development of CRQC, the mathematical problems underpinning modern security—such as integer factorization and discrete logarithms—are no longer computationally hard. This framework quantifies the urgency of migration to prevent systemic collapse.
This research introduces a coordinate-based analytical framework to study the co-evolution of two distinct technological trajectories: the evolution of cryptography (from pre-quantum to quantum-safe) and the evolution of computing (from classical to Cryptographically Relevant Quantum Computing). The evolution is defined by a transition from static, classical security models to dynamic, quantum-resilient infrastructures. We track this via the interplay of cryptographic standards and computational power, mapping the trajectory from the legacy era to the post-quantum equilibrium.
@@ II Overview of the Framework
III Acceleration of Quantum Threats
III-A Algorithmic Optimization for Cryptanalysis
III-B Hardware-Specific Design Optimization
III-C Algorithmic-Hardware Co-optimization
III-D Differential Analysis of Classical Hardness
IV Phases of Quantum-Cryptographic Evolution
IV-A Quadrant (-, -): Classical Pre-Quantum Era
@@
  • Classical pre-quantum systems: These systems rely on RSA and ECC primitives. They are standard in non-CRQC environments and assume that integer factorization remains computationally hard for the foreseeable future.
  • Prequantum cryptography: Algorithms are based on integer factorization and discrete logarithms. These primitives have served as the backbone of global trust for decades, providing the foundation for secure internet communications.
@@
  • Rudimentary Quantum computing: NISQ-era devices exist but lack the depth for Shor’s algorithm. They represent the infancy of quantum hardware, focusing on experimental physics rather than large-scale computation.
@@
  • PQC and Hybridization: Initial PQC standards are announced. Research identifies gaps in early PQC implementations, necessitating the development of hybrid cryptographic models to ensure backward compatibility.
@@
IV-B Quadrant (+, -): Proactive Transition
@@
IV-C Quadrant (-, +): Vulnerability Crisis
@@
IV-D Quadrant (+, +): Dynamic Equilibrium
@@
V Background: Quantum-Cryptographic Evolution
@@ V-C Quantum Networks and Their Evolution
V-D Evolution of Quantum Threats
V-D1 Threats to Public-Key Cryptosystems
V-D2 Threats to Symmetric-Key and Hashing
V-D3 Harvest Now, Decrypt Later (HNDL)
V-E Evolution of Cryptographic Standards
@@
V-F Evolution of Mathematical Assumptions and Hardness
@@
V-G Evolution of Quantum Resilience
@@
VI Discussions
The coordinate system reveals that security is not a static destination but a moving target. The (,+)(-,+) quadrant represents a point of no return for data confidentiality. The transition from (,)(-,-) to $(+,+)$ must be handled with extreme focus on crypto-agility, as the (+,)(+,-) phase is the only viable path to avoid the crisis of (,+)(-,+). The emergence of novel attacks in the $(+,+)$ quadrant suggests that cryptographic resilience will require perpetual innovation and a shift from project-based security to operational resilience.
Figure 2 presents the overall four quadrants.
VI-A Future Trajectory and Persistent Challenges
@@ VII Related Work
VIII Conclusions
The evolution of cryptography is an arms race. Organizations must move toward the $(+,+)$ equilibrium by institutionalizing crypto-agility and proactive PQC adoption. Failure to do so results in permanent loss of data confidentiality via HNDL.

### 22 Apr 2026 v2

@@ Abstract
Quantum computing poses an existential threat to global cryptographic infrastructure. We study the evolution of quantum vs cryptography, and the systemic risk posed by Cryptographically Relevant Quantum Computers (CRQC). We propose a two-dimensional coordinate system to map the co-evolution of cryptographic resilience and computational capability. Additionally, we introduce a hyper-dimensional framework to taxonomize evolutionary pressures across seven fundamental dimensions. Our contributions are: (1) a coordinate-based risk framework, (2) a seven-dimensional taxonomy of quantum-cryptographic evolution, (3) an analysis of threat acceleration factors, and (4) a roadmap for quantum-resilient migration. This framework should be used by organizations and systems for risk assessment, infrastructure planning, and managing the transition to crypto-agile architectures.
@@ I Introduction
The emergence of CRQC renders the mathematical foundations of modern security—such as integer factorization and discrete logarithms—obsolete. This creates a ”Quantum Gap,” representing the systemic risk between the arrival of quantum threats and the deployment of quantum-safe defenses. This research introduces a coordinate-based analytical framework to study the co-evolution of two distinct technological trajectories: the evolution of cryptography (from pre-quantum to quantum-safe) and the evolution of computing (from classical to quantum). We track this via the interplay of cryptographic standards and computational power, mapping the trajectory from the legacy era to the post-quantum equilibrium.
I-A Contributions
This paper makes the following primary contributions:
  1. 1.
    Coordinate-Based Framework: We propose a 2D quadrant model to visualize systemic risk and the urgency of cryptographic migration.
  2. 2.
    Hyper-Dimensional Taxonomy: We introduce a seven-dimensional framework (D1-D7) that categorizes the co-evolution of quantum algorithms, hardware, networks, and threat vectors.
  3. 3.
    Acceleration Analysis: We quantify the factors—ranging from T-gate reduction to hardware co-design—that are accelerating the timeline for CRQC-driven cryptanalysis.
  4. 4.
    Quantum-Resistance Roadmap: We provide a comprehensive analysis of the phases of quantum-cryptographic evolution, offering a path for systems to move from legacy states to dynamic quantum-resilience.
I-B Organization of the Paper
The remainder of this paper is organized as follows: Section II presents the overall framework to express the quantum-cryptographic evolution. Section III details the phases of evolution across the four quadrants. Section IV presents the hyper-dimensional framework for quantum-cryptographic co-evolution. Section V analyzes the evolution. Section VI presents the evolution of quantum threats. Section VII analyzes the factors accelerating quantum threats. Section VIII and Section IX present evolution of quantum resilience and implementation and certification of quantum resilient cryptographic techniques such as PQC standards. Section X discusses the future trajectory and challenges. Related work is reviewed in Section XI, and Section XII concludes the paper.
@@ II Overview of the Framework
III Phases of Quantum-Cryptographic Evolution
III-A Quadrant (-, -): Classical Pre-Quantum Era
@@
  • Pre-quantum cryptography: Classical cryptography primitives not designed for to be quantum-resistant. These primities and cryptosystems may not be quantum resistant or maybe quantum resistant. Primitives based on integer factorization and discrete logarithms are quantum-unsafe. These primitives have served as the backbone of global trust for decades, providing the foundation for secure internet communications. However, AES-256 even though pre-quantum is considered quantum-safe at least for now.
  • Classical pre-quantum systems: These systems rely on RSA and ECC primitives. They are standard in non-CRQC environments and assume that integer factorization remains computationally hard for classical computing for the foreseeable future.
  • Basic Quantum computing: NISQ-era devices exist but lack the depth for Shor’s algorithm. They represent the infancy of quantum hardware, focusing on experimental physics rather than large-scale computation.
@@
  • Post-quantum cryptography (PQC): NIST announced post-quantum cryptography standards in the year of $2024$ after about eight years of efforts. The PQC standardization is in in response to protection against quantum threats such as Shor’s algorithm and Grover’s algorithm. Research identifies gaps in early PQC implementations, necessitating the development of hybrid cryptographic models to ensure backward compatibility.
  • Hybrid Cryptography: Research identifies gaps in early PQC algorithms, implementations, necessitating the development of hybrid cryptographic models to ensure stronger quantum-resistance and backward compatibility.
@@
III-B Quadrant (+, -): Proactive Transition
@@
III-C Quadrant (-, +): Vulnerability Crisis
@@
III-D Quadrant (+, +): Dynamic Equilibrium
@@
III-E Quantum Vs Cryptography Quadrants
The coordinate system reveals that security is not a static destination but a moving target. The (,+)(-,+) quadrant represents a point of no return for data confidentiality. The transition from (,)(-,-) to $(+,+)$ must be handled with focus on PQC and quantum-resistant approaches, their adoption alongwith crypto-agility, as the (+,)(+,-) phase is the only viable path to avoid the crisis of (,+)(-,+). The emergence of novel attacks in the $(+,+)$ quadrant suggests that cryptographic resilience will require perpetual innovation and a shift from project-based security to operational resilience. Figure 2 presents the overall four quadrants.
IV A Hyper-Dimensional Framework for Quantum-Cryptographic Co-evolution
The evolution of quantum-cryptographic systems is characterized by a complex interplay of multiple dimensions. These dimensions evolve both independently and within interconnected subgroups, where progress in one plane frequently influences the trajectory of another. This co-evolutionary process ultimately manifests as the emergence of quantum threats to existing cryptographic standards. To organize, study, and analyze this evolution, we propose a taxonomy based on the following hyper-dimensional framework.
IV-A Taxonomy of Dimensions
The framework is defined by the following dimensional systems:
  • D1 (Foundational Dimensions): Comprises (1) Information theory, (2) Computational complexity, and (3) Mathematics.
  • D2 (Information Modality): Defines the referential frame for information processing: (1) Quantum and (2) Classical.
  • D3 (Operational Dimensions): Represents the three primary domains of technological application: (1) Computing, (2) Network.
  • D4 (Developmental Stage): Distinguishes between the (1) Theoretical and (2) Practical stages of development.
  • D5 (Cryptographic Core): Represents the primary domains of (1) Cryptography and (2) Algorithms.
  • D6 (Implementation Layer): Influenced by D5, this system includes (1) Algorithms to solve mathematical problems of cryptography, algorithms to search (2) Theoretical implementation, (3) Practical implementation of algorithms.
  • D7 (Quantum Threats): The resultant dimension representing the evolution of threats, including (1) Harvest Now, Decrypt Later (HNDL), (2) Compromise of public-key cryptography, (3) Compromise of symmetric-key cryptography, and (4) Compromise of other cryptographic primitives and protocols.
TABLE VI: Hyper-Dimensional Framework Taxonomy
Dimension Sub-dimensions Role
D1 Info Theory, Comp. Complexity, Math Foundational
D2 Quantum, Classical Modality
D3 Computing, Memory, Network Operational
D4 Theory, Practice Developmental
D5 Cryptography, Algorithms Core
D6 Math Solutions, Theory Impl, Practical Impl Implementation
D7 HNDL, PKI Compromise, Sym. Compromise, Other Threat Outcome
D1: FoundationalD2: ModalityD3: OperationalD4: DevelopmentalD5: CoreD6: ImplementationD7: Quantum Threats
Figure 3: Evolutionary Influence Graph: The flow of evolutionary pressure across dimensions.
Figure 3 provides a macro-perspective of the framework, mapping the primary evolutionary pressure flow from foundational dimensions (D1) to the resultant quantum threats (D7). The solid arrows represent the primary causal chain of development, while dashed lines denote secondary cross-dimensional dependencies, illustrating how progress in foundational theory (D1) or operational domains (D3) can directly catalyze changes in the implementation layer (D6) and the cryptographic core (D5).
IV-B Influence Graph
Figure 4 maps the dense inter-dimensional dependencies.
D1.1D1.2D1.3 D1D2.1D2.2 D2D3.1D3.2D3.3 D3D4.1D4.2 D4D5.1D5.2 D5D6.1D6.2D6.3 D6D7.1D7.2D7.3D7.4 D7
Figure 4: Detailed Representative Influence Graph: Some inter-dimensional dependencies are depicted.
Figure 4 offers a granular decomposition of this framework, mapping specific sub-dimensions (e.g., D1.1, D1.2) within their respective hierarchical groups. The graph maps the dense inter-dimensional dependencies where each edge represents a directed evolutionary pressure. For instance, advancements in D1.2 (Computational Complexity) directly constrain the design of D5.2 (Algorithms), while D3.1 (Quantum Computing) accelerates D6.3 (Practical Implementation).
This visualization allows researchers to track evolutionary progress at the component level, where specific sub-dimensions undergo independent maturation, providing a clear view of how individual advancements aggregate to form the higher-level dimensions of the co-evolutionary system.
The interoperation across these dimensions suggests that the evolution of quantum threats is not a linear progression but a multi-variate outcome of advancements across the foundational, operational, and implementation layers of the hyper-dimensional framework.
V Analysis of the Evolution
@@ V-C Quantum Networks and Their Evolution
V-D Evolution of Cryptographic Standards
@@
V-E Evolution of Mathematical Assumptions and Hardness
@@
VI Evolution of Quantum Threats
VI-1 Threats to Public-Key Cryptosystems
VI-2 Threats to Symmetric-Key and Hashing
VI-3 Harvest Now, Decrypt Later (HNDL)
VII Acceleration of Quantum Threats
TABLE VII: Factors Accelerating Quantum Cryptanalytic Threats
Factor One-Line Summary
Algorithmic Refinement Reduction in T-gate counts and circuit depth for Shor’s algorithm.
Hardware Topology Transition to reconfigurable neutral atom arrays for dynamic gate execution.
Error Correction Deployment of LDPC and surface codes to suppress logical qubit decoherence.
Co-design Synergy Direct mapping of optimized quantum circuits onto specialized hardware architectures.
Hardness Analysis Prioritization of ECC-based assets due to lower qubit requirements for breakage.
Resource Scaling Recent breakthroughs reducing logical qubit requirements for RSA-2048 to 100,000.
VII-A Algorithmic Optimization for Cryptanalysis
VII-B Hardware-Specific Design Optimization
VII-C Algorithmic-Hardware Co-optimization
VII-D Differential Analysis of Classical Hardness
VIII Evolution of Quantum Resilience
@@
IX Evolution of Implementation Security and Certification
The implementation of quantum-safe and post-quantum cryptographic algorithms and protocols represents a critical dimension in the evolution toward quantum resilience. Beyond the mathematical robustness of the primitives, the security of these systems is fundamentally contingent upon the rigor of their implementation, testing, and hardening processes. While these implementations aim to mitigate quantum threats, they simultaneously introduce a potential expansion of the attack surface, ranging from classical side-channel vulnerabilities to novel quantum-assisted implementation attacks. Consequently, established security standards such as FIPS 140-3 [16] and NIST SP 800-140 [38] must undergo significant evolution to address the unique requirements of PQC, ensuring the correctness and security of both software and hardware modules. Furthermore, cryptographic certification frameworks, including ISO/IEC standards [39], must transition to accommodate these new primitives, providing a verifiable path to ensure that quantum-resistant deployments do not inadvertently introduce systemic weaknesses. This evolution of certification is a prerequisite for moving from the proactive transition phase to a state of dynamic equilibrium.
X Future Trajectory and Challenges
The coordinate system reveals that security is not a static destination but a moving target. The (,+)(-,+) quadrant represents a point of no return for data confidentiality. The transition from (,)(-,-) to $(+,+)$ must be handled with extreme focus on crypto-agility.
@@ XI Related Work
XII Conclusions
The evolution of cryptography is an arms race. Organizations must move toward the $(+,+)$ equilibrium by institutionalizing crypto-agility and proactive PQC adoption. We demonstrate that the co-evolution of quantum threats and cryptographic resilience requires a holistic, hyper-dimensional approach to security, moving beyond static implementation toward a framework of perpetual monitoring and rapid algorithmic renewal. Failure to institutionalize these practices results in permanent loss of data confidentiality via HNDL. Future challenges include the emergence of quantum-assisted side-channel attacks and the need for ”Continuous Cryptographic Agility.”